OpenStudy (anonymous):

Which two square roots are used to estimate the square root of 42

TheSmartOne (thesmartone):

Welcome to OpenStudy!

TheSmartOne (thesmartone):

so we need to find out what square root of 42 is in between so list all the squares of number can you tell me what these equal to: 2^2 = ? 3^2 = ? 4^2 = ? 5^2 = ? 6^2 = ? 7^2 = ? 8^2 = ?

TheSmartOne (thesmartone):

I'll do a few for you so you see how to do it 2^2 = 2*2 = 4 3^2 = 3*3 = 9 4^2 = 4*4 = 16 can you do the rest? @Mathiscoolitsgreat

OpenStudy (anonymous):

25 36 49 64

TheSmartOne (thesmartone):

ok, and in between which two of those numbers is 42 in? choose one number which is the next number larger than 42, and choose another number which is the next number that is smaller than 42 for example, 36 is the next number after 28, and 25 is the number before 28

TheSmartOne (thesmartone):

in these three numbers 36 49 64 can you tell me which two numbers that 42 is in between?

OpenStudy (anonymous):

36 and 49 ?

TheSmartOne (thesmartone):

correct, and so we can find that the square root of 42 has to be in between the square root of 49 and square root of 36 \(\sf\Large \sqrt{36} < \sqrt{42} < \sqrt{49}\) and can you tell me whether 42 is closer to 49 or closer to 36?

OpenStudy (anonymous):

36

TheSmartOne (thesmartone):

correct and since 42 is practically in between sqrt of 36 and sqrt of 49 but it is a little closer to sqrt of 36 we can estimate that square root fo 42 has to be smaller than 6.5 since it's closer to 6 than it is to 7 note: square root of 36 = 6 square root of 49 = 7

TheSmartOne (thesmartone):

and sqrt of 42 is approximately 6.48 so yay!

OpenStudy (anonymous):

Thank you for the help!!

TheSmartOne (thesmartone):

Anytime!

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8